# Dynamical mass of a solar-like oscillator at the main-sequence turnoff from Gaia astrometry & ground-based spectroscopy

### v1

@@ A.1 Fundamental parameters
The code also automatically adjusts a mean line-broadening value to fit the Fe line profiles. The projected rotational velocity () was then estimated by assuming that this broadening is the quadratic sum of the instrumental resolution (), macroturbulence (adopted from the prescription of Doyle et al. (2014) for G–K main-sequence stars), and rotational broadening. The derived fundamental parameters are reported in Table 2.
@@ B.4 Individual frequency modeling with MESA
Figure 8 shows the position of KIC 9693187 according to the spectroscopic and seismic surface gravity in blue and purple, respectively. Furthermore we mark the position of the best fit model from the individual frequency modeling in the HRD. To test if KIC 9693187’s primary is located before or after the turnoff from the mainsequence, we also marked the depletion of the fractional content of hydrogen in the core, whereby we mark the positions X and , confirming that the star has consumed almost its core hydrogen.

### v2

@@ A.1 Fundamental parameters
The code also automatically adjusts a mean line-broadening value to fit the Fe line profiles. The projected rotational velocity () was then estimated by assuming that this broadening is the quadratic sum of the instrumental resolution (), macroturbulence (adopted from the prescription of Doyle et al. (2014) for G–K MS stars), and rotational broadening. The derived fundamental parameters are reported in Table 2.
@@ B.4 Individual frequency modeling with MESA
Figure 8 shows the position of KIC 9693187 according to the spectroscopic and seismic surface gravity in blue and purple, respectively. Furthermore we mark the position of the best fit model from the individual frequency modeling in the HRD. To test if KIC 9693187’s primary is located before or after the turnoff from the MS, we also marked the depletion of the fractional content of hydrogen in the core, whereby we mark the positions X and , confirming that the star has consumed almost its core hydrogen.